OpenStudy (anonymous):

how do I go about determing the polynomial equation that passes through these 4 points? (-3,10), (-1,0), (0,-5), (3,52)

OpenStudy (swissgirl):

Hmmm are you familiar with matrices? What method should you be using?

OpenStudy (anonymous):

it doesn't matter

OpenStudy (swissgirl):

Ok what grade r u in?

OpenStudy (anonymous):

11

OpenStudy (swissgirl):

hmmm ok we will try matrices but I doubt u learnt that method. It may be too advanced

OpenStudy (anonymous):

I know a little about matrices, mainly how to set them up using the coefficients but that's all

OpenStudy (swissgirl):

Ok for starters we have points so we are assuming that our polynomial is of degree 3 ok?

OpenStudy (anonymous):

right

OpenStudy (swissgirl):

Now we know that a polynomial of degree 3 has the form of \(f(x)=a+bx+cx^2+dx^3\)

OpenStudy (anonymous):

ok-that's backwards from how I normally write it but ok

OpenStudy (swissgirl):

No we were given 4 diff points so we will plug them in Our first point is (-3,10) \(a+b(-3)+c(-3)^2+d(-3)^3=10\) \(a-3b+9c-27d=10\) Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do that with each set of points and then what

OpenStudy (swissgirl):

Ok so our next point is (-1,0) \(a+(-1)b+(-1)^2c+(-1)^3d=0\) \(a-b+c-d=0\) Our next point is (0,-5) \(a+(0)b+(0)^2c+(0)^3d=-5\) \(a=-5\) Our next point is (3,52) \(a+(3)b+(3)^2c+(3)^3d=52\) \(a+3b+9b+27d=52\)

OpenStudy (swissgirl):

Ok now that we know that a=-5 we can plug that into all the 3 equations

OpenStudy (anonymous):

ok which equation is the best one to use

OpenStudy (swissgirl):

No you gotta plug a=-5 into all the 3 equations

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i have done that with all 3 now what

OpenStudy (swissgirl):

Equation #1: \((-5)-3b+9c-27d=10\) \(-3b+9c-27d=15\) Equation #2 \((-5)-b+c-d=0\) \(-b+c-d=5\) Equation #3 \((-5)+3b+9c+27d=52\) \(3b+9c+27d=57\)

OpenStudy (anonymous):

that's what i got

OpenStudy (swissgirl):

Ok next we will set up everything into a matrix The first column will be our b's second will be c's and third will be d's and the last column will be our answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

not sure what goes in the last column

OpenStudy (swissgirl):

\[ \left( \begin{array}{cccc} -3 & 9 & -27 &15 \\ -1 & 1 & -1 &5\\ 3 & 9 & 27 &57 \end{array} \right)\]

OpenStudy (swissgirl):

This is how our matrix shld look

OpenStudy (anonymous):

good so far

OpenStudy (swissgirl):

Now we need to row reduce which is a pain Do you know how to row reduce?

OpenStudy (anonymous):

no

OpenStudy (swissgirl):

ughhhh Shi*

OpenStudy (anonymous):

sorry

OpenStudy (swissgirl):

Well it will take me a few minutes to show you each step. Its really hard to show that matrix thing here. Ill try to explaing as much as possible

OpenStudy (swissgirl):

Ok first I will switch the rows \(\left( \begin{array}{cccc} -1 & 1 & -1 &5\\ -3 & 9 & -27 &15 \\ 3 & 9 & 27 &57 \end{array} \right)\)

OpenStudy (swissgirl):

Then I multiplied the first row by -1 So -1*Row #1 \(\left( \begin{array}{cccc} 1 & -1 & 1 &-5\\ -3 & 9 & -27 &15 \\ 3 & 9 & 27 &57 \end{array} \right)\)

OpenStudy (swissgirl):

Next I Multiply Row #1by 3 and add that to Row#2 Row#2=3*Row#1+Row#2 \(\left( \begin{array}{cccc} 1 & -1 & 1 &-5\\ 0 & 6 & -24 &0 \\ 3 & 9 & 27 &57 \end{array} \right)\)

OpenStudy (anonymous):

why mult. by 3

OpenStudy (swissgirl):

Next you divide Row#2 by 6 Row#2=Row#2/6 \(\left( \begin{array}{cccc} 1 & -1 & 1 &-5\\ 0 & 1 & -4 &0 \\ 3 & 9 & 27 &57 \end{array} \right)\)

OpenStudy (swissgirl):

OK basically we are trying to get the matrix into this form \(\left( \begin{array}{cccc} 1 & a & b &c\\ 0 & 1 & d &e \\ 0 & 0 & 1 &f \end{array} \right)\) Where a,b,c,d,e,f are numbers I just giving you an example. Just trying to give u an idea So basically work row by row We first get the first number in row 1 to be 1 Then we go to the next row and get the first number to be 0 and the second one to 1 And then we go to the third row and get the first 2 numbers to be 0 and the third one to be 1

OpenStudy (anonymous):

not sure what to do to get equation

OpenStudy (swissgirl):

OK so basically lets look at the multiplying 3 one that you got confused by ok

OpenStudy (swissgirl):

In the second Row the first number is -3 correct?

OpenStudy (anonymous):

how did u get -3

OpenStudy (swissgirl):

\(\left( \begin{array}{cccc} 1 & -1 & 1 &-5\\ -3 & 9 & -27 &15 \\ 3 & 9 & 27 &57 \end{array} \right)\) This our matrix we had in the beg

OpenStudy (anonymous):

yes way back up at the top

OpenStudy (swissgirl):

Correct So u see that -3 ? Its the first number in the second row

OpenStudy (anonymous):

yes

OpenStudy (swissgirl):

Ok we need it to be 0 correct?

OpenStudy (anonymous):

why

OpenStudy (swissgirl):

Cuz we need it to be in that form \(\left( \begin{array}{cccc} 1 & a & b &c\\ 0 & 1 & d &e \\ 0 & 0 & 1 &f \end{array} \right)\)

OpenStudy (anonymous):

k

OpenStudy (swissgirl):

Ok so how can we get the -3 to become a 0?

OpenStudy (anonymous):

add 3

OpenStudy (swissgirl):

Well we can't do that Cuz remember the whole row in one equation remember? And there are rules to matrices Firstly you can only divide and multiply a single row And secondly you can add and subtract rows

OpenStudy (swissgirl):

You cant just add -3 to a whole row So what I did was multiply Row#1 by 3 and then add that to Row #2 So our new Row#2 =3*Row#1+Row#2

OpenStudy (swissgirl):

Like I think this method is too advanced for you

OpenStudy (anonymous):

i agree

OpenStudy (swissgirl):

hmmmm thinking how else to solve this give me a sec

OpenStudy (swissgirl):

hahahha ok So I have another way to solve Using elimination Are you familiar with that? I remember using that method back in HS so i bet u do

OpenStudy (anonymous):

I have used it before

OpenStudy (swissgirl):

Ok so lets write out the 3 equations again so it will be easier

OpenStudy (swissgirl):

-3b+9c-27d=15 -b+c-d=5 3b+9c+27d=57

OpenStudy (anonymous):

I have c+4

OpenStudy (anonymous):

=4

OpenStudy (swissgirl):

Ok so we will set subtract equation #1 and equation #2 And we will subtract Equation #1 and equation #3 We will eliminate the "b" ok And then we will have 2 brand new equations \(-3b+9c-27d=15\) \(-3( -b+c-d=5))\) We can only eliminate the b by multiplying the second equation by -3 \(-3b+9c-27d=15\) \(3b-3b+3b=-15\) --------------------- 0b+6c-24b=0 ----> Equation #4 \(-3b+9c-27d=15\) \(3b+9c+27d=57\) --------------------- 0b+18c+0d= 72 -----> Equation #5 So 18c=72 \(\large {18c \over 18} ={72 \over 18}\) c=18 Ok so now lets plug in c=18 into equation #4 6(18)-24b=0 108-24b=0 -24b=-108 b=4.5 Now lets plug c=4 and b=4.5 into Equation #2 (-4.5)+4-d=0 -.5-d=0 -.5=d

OpenStudy (anonymous):

my choicces are x^3+4x^3+2x-5 or x^3-4x^2-2x-5

OpenStudy (swissgirl):

hmmm we got the wrong answer let me recheck my work

OpenStudy (anonymous):

and we came up with a positive fouand a -5 but not a 1 or a 2

OpenStudy (anonymous):

positive four

OpenStudy (swissgirl):

ohh I see where i went wrong ugghhh

OpenStudy (anonymous):

u used 18 instead of 4 for c

OpenStudy (swissgirl):

yesssssssssssssssss

OpenStudy (swissgirl):

Ok so now lets plug in c=4 into equation #4 6(4)-24b=0 24-24b=0 -24b=-24 b=1 Now lets plug c=4 and b=1 into Equation #2 -(1)+4-d=0 3-d=0 3=d

OpenStudy (swissgirl):

Still getting a diff answer

OpenStudy (swissgirl):

hmmmmm

OpenStudy (anonymous):

this might be a problem that needs to be done on a calculator. I think I will just mark the first answer listed with the positive four. I didn't expect all of this help from you. thanks so m;uch

OpenStudy (swissgirl):

Ill brb I just need to eat dinner. Everything is cold by now

OpenStudy (anonymous):

I see, First off, I'd graph it.

OpenStudy (anonymous):

but how do I get an equation from a graph

OpenStudy (swissgirl):

Nah u r not suppossed to graph this stuff. Its 3 variable so it will be annoying

OpenStudy (swissgirl):

ok Let me go over my work. and see where i messed up

OpenStudy (anonymous):

I have a few ways although I'm just checking if they work first.

OpenStudy (swissgirl):

Ohhh I see where i messed up lol

OpenStudy (swissgirl):

hahah I totally mixed up variables all over

OpenStudy (swissgirl):

Today isnt my day, by accident i wrote out a check for $300 instead of $800 I think its the lack of food OK let me start form way back there

OpenStudy (anonymous):

f(x) =af(k(x-d))+c It was -5 down from the y-axis axis so the c is is -5

OpenStudy (anonymous):

Lagging so much it's unbelievable.

OpenStudy (anonymous):

the d is 0 since the x-axis is the same.

OpenStudy (anonymous):

you mean d is -5 ax^4+bx^3+cx^2+x

OpenStudy (swissgirl):

−3b+9c−27d=15 3b−3c+3d=−15 --------------------- 0b+6c-24d=0 ----> Equation #4 −3b+9c−27d=15 3b+9c+27d=57 --------------------- 0b+18c+0d= 72 -----> Equation #5 18c=72 c=4 Now we plug that c=4 into equation #4 6(4)-24d=0 24-24d=0 -24d=-24 d=1

OpenStudy (anonymous):

my choices listed all have d as -5

OpenStudy (anonymous):

Oh wow Im lagging wayy too much . Sorry.

OpenStudy (anonymous):

right swiss, your d shoulld be one

OpenStudy (swissgirl):

No ure d=1 in all cases Remember the formula f(x)=a+bx+cx^2+dx^3

OpenStudy (anonymous):

d - 1 in all my choices if written your way

OpenStudy (anonymous):

a= -5 in all my choices

OpenStudy (anonymous):

c is either 4 or -4 and b is either 2 or -2

OpenStudy (swissgirl):

a=-5 as we saw in the beg c=4 d=1 and now we gotta show what b is

OpenStudy (anonymous):

-2